- Status Report
- Feb 8, 2023
ISS MER Status Report 15 Dec 2000
Note: the following is contained within “Felicity Vol 040” ISS Integrated Program Schedule, 15 December 2000 (1.7 MB Adobe Acrobat file)
ISS MER Status Report
1- Shortly after undocking, a limit switch on the Solar Array Blanket Box was indicating latched
and unlatched. The sensor has continued to toggle between the two positions during day and
night. Indications are that the sensor is toggling due to the thermal cycles. ISS MER is still
2- FGB Battery 1 completed Capacity Restoration Mode on 12/14. Battery 1 is currently in
3- Crew performed trouble shooting procedure on the STBD UOP on 12/13. The UOP tripped
the RPC during the test and is considered failed. Future trouble shooting will be performed on
the “failed” UOP.
4- Port UOP #2 tripped several times on 12/13 while the crew was using the SSC computer.
The crew performed the troubleshooting procedure on the port UOP and the UOP worked
nominally. The crew also connected the SSC and could not recreate the trip. Efforts are still
underway to troubleshoot other causes as identified on the Fault Tree.
At GMT 250/09: 22: 28 Battery #6 was commanded to go into the Cycle mode. After Battery #6 entered the
Cycling Mode, Battery #1 also entered Cycling Mode. The Cycling Mode termination command was issued
via ECS and a second set of commands to initiate Cycling Mode of Battery #6 was uplinked via ECS. Again
Battery #1 entered Cycling together with Battery #6. After DO3 comm pass, Battery #6 ZRU was
disconnected from the Main Bus via ECS. This issue is under investigation by the Russians. However both
battery 6 &1 are presently on line with nominal performance.
At GMT 258: 04: 38, during a Russian Ground site, it was noted that SM Battery 6 was in cycle mode
unexpectedly. RSC- E confirmed that this battery was not commanded into cycle mode. Investigation of the
KhSC commanding is currently in work to determine if an erroneous command sent from KhSC could have
cause this signature. SM Battery 6 is currently on line and performing nominal.
For both PCU 1 & 2, the Tube Temperatures are 20 deg C below their expected value. The tube temperatures should have
reached operating temperature and started cycling between 32 and 42 deg C, but have steadied out at approximately 22 & 25 deg
C each. The Tube Heaters remain ON constantly, providing 15W of heat to the tubes. Two possibilities are being investigated. 1)
The Tube Heater wattage is too small or 2) The Tube Temperature Sensor is reading the wrong temperature.
At GMT 2000/333: 14: 49: 03. The Station crew reported that the CKV Fan 2 either disintegrated or ingested an object that had the
effect of disintegrating one or both of the CKV 2 fans. If only one fan is affected, the flow is reduced, but capability remains. If
both fans are affected, it basically means we have lost SVK 2 (air conditioner 2). The crew reported that a blade broke off the fan
and moved to an unknown location. As of GMT 2000/333: 20: 18, it is believed the other fan for SVK 2 is OK; but will not operate
until blade and other debris are retrieved.
During the Loop B Z1/P6 QD mating it was noted that the Trailing and Starboard Radiator accumulator quantity sensors for Loop A
changed. And when the Loop A QD was mated it was noted that the Trailing and Starboard Radiator accumulator quantity
sensors for Loop B changed. Upon inspection of data taken during PFCS activation and during EEATCS Loop Activation on GMT
341/00: 35, it was determined that the most likely cause is cross- wiring of the radiator accumulator sensors.
Temperature sensor data indicates that one of the BCDU base plate temperature sensors has an incorrect
gain coefficient that makes it deviate from the family of other base plate temperature sensors. Data
retrieved from the archives of the IACO test shows that the deviation was present in that test environment.
A caution will be generated if the temperature exceeds the caution limit. Effort is on the way to determine a
new gain coefficient or offset for the temperature sensor PPL.
Several battery heater trips occurred during the initial activation procedure and have not reoccurred since
then. These trips occurred when the battery heater power switch in the BCDU was closed and the heater
switch in the battery was also closed. The event was encountered during MEIT testing and PRACA 1483
was opened. The system exposure to the trip of the heater switch is limited to the first few orbits while the
initial charge of the batteries is being conducted. This anomaly will only occur when the heater switch is
closed and the battery heater switch has already been closed. The trip can be reset immediately after the
event. Documented in Flight Rule B9. 5. 10- 11. Issue is still under investigation.
When the launch restraints for the 2 four- bar mechanisms for the Beta Gimbal Assemblies were removed, spring forces should
have initiated motion for the assemblies. Instead, the mechanisms didn’t move and the crew had to deploy them manually. The
springs were expected to initiate motion, with crew required to manually complete deployment. There was uncertainty as to
exactly how much the springs would accomplish and how much would be left for the crew. As it happened, the springs didn’t
contribute as much as was expected, requiring full manual deployment. This issue is under investigation.
After PVCU MDM activation at GMT 338/22: 20, the EEATCS loop A & B temps were both reading around -9 deg C. At these
temperatures, it would be expected that the low temp. message would be issued once and remain in- alarm. Currently the
EEATCS Loop A& B Outlet temps are within limits. A test is scheduled for 12/18 to test the Min Out FDIR response.
crew reported A/C1 was noisy and vibrating since activation. Energia indicated they suspect A/C1 power supply unit is faulty due
to a low voltage reading. A/C1 was shutdown on GMT 341and faulty A/C1 power supply unit was down manifested by 4A for
Crew reported that A/C2 fan vanes sheared off and one stator was bent when it was “inadvertently” activated on GMT 331. A/C2
fan was R& R’ed and A/C2 was activated on GMT 341. We have lost redundancy in air conditioning function since we now have
no spare power supply unit on board. Energia has requested we launch a spare air conditioning power supply unit on 5A.
“J3” connector label came loose from Z1 Patch Panel Z1- 006. Issue is under investigation
Node 1 SVS target on end cone panel NOD1C208 had large scratch on black inner circle. Anomaly noted
on STS- 96 (2A. 1). EBIT developed SVS workaround plan, in place for future SRMS operations.
RPC feeding UOP 1 tripped due to over current. UOP had two loads connected to it – FPP NCU and
KuBand power supply. Actual RPC output over current was confirmed with data dump. Initial trouble
shooting saw a reoccurrence of the trip with no loads attached to the UOP. UOP has failed, and Initial
trouble shooting has determined that it should be red tagged (not used). Further trouble shooting plans are
SABB Latch indication are intermittently being lost. Problem first noticed during 4A undocking. Data
investigation shows that it was occurring previous to this event. The issue is under investigation and a
PRACA is being generated.
System (BRS) Pin failed to release on 1st attempt
BRS Pin is a spring loaded device, however one unit failed to release upon initial actuation. P6 restored on-
orbit via second actuation attempt. Known potential problem prior to flight (tolerance analysis identified
potential for failure to actuate), EVA tool had been designed and manifested in case of complete failure to
Device (RTD) Worksite
During EVA #1 and EVA #2 (GMT 340/22: 30: 00) and EVA# 3 (GMT 342/20: 26: 00), the EVA Crew removed
the radiator Resistive Thermal Devices (RTDs) and several pieces of Kapton tape, that was used to anchor
the RTD lead wires to structure, came loose from the wire and were lost. It appeared that there was
excessive Kapton tape used to anchor the RTD lead wires to the radiator structure. BCP Kapton Tape wire
anchor procedure will be reviewed and updated to insure that the tape anchor is required and if used, to be
secured to the wire and that a maximum tape anchor spacing be specified and used during installation.
Batt 2B22 Thermal Data Filter Pressure Sensor Failure- P6 advisory was seen to go in alarm and then back
to normal nine times. This advisory indicates that one of the Battery cell pressure sensors had deviated
more than 30 psi from the average pressure measurement and that it had been taken out of the algorithm
which calculates Battery State of Charge (SOC). Sensor offset on this Battery was seen in MEIT and IACO
testing and as a result PPL 18 version 102 was built. No impact to the present Battery operation. However,
if another cell pressure sensor deviates by 30 psi, the SOC algorithm will stop working and the SOC will
remain at the last calculated value and SOC invalid parameter will be observed. Battery charging will occur
at the rate specified by the invalid SOC and will not decrease as the battery charges.
Solutions being considered are:
1- Halt charging of that Battery set if the SOC invalid parameter is observed.
2- Evaluate downloading PPL 18 to widen the range for declaring a pressure sensor failed.
evaluating a new PPL to apply a scaling factor to correct the cell pressure sensor error.
3- Modify the gain & offset PPL for the battery to correct the cell pressure offset.
UOP 2 GFCI tripped while the SSC laptop was connected and in use. GFCI was reset by the crew but it tripped several times
again. SSC laptop, power supply and cable are suspect. Further trouble shooting is planned.
The crew has not been able to set up the SSC network. They have tried several trouble shooting plans without satisfactory
results. A temporary work around is used to provide temporary connection. The issue is under investigation.
During the 2P Progress docking to the ISS FGB nadir port, the Progress Kurs system was unable to “lock on” to the FGB (- Y)
directional antennae, as needed for the final berthing sequence. The automated rendezvous and docking proceeded nominally
until the station- keeping point was reached. The in- coming Progress acquired and tracked the SM- omni antennae and performed
a nominal fly around. However, when the final berthing sequence from the station- keeping point began, and the Progress tried to
switch from the SM omni antennae to the directional antenna located near the docking port [approximate GMT 323: 02: 55], a
consistent data link from the FGB “Sector Acquisition” and “Scanning Orientation” antennae was never achieved. These
antennae provide range, range rate, and relative attitude information for the final berthing sequence. At approximately 323: 03: 05,
manual control of the docking vehicle was seized by the ISS crew via the TORU system, and the Progress was successfully
docked. The issue is under investigation.